OpenStudy (anonymous):

You roll a standard number cube. Find P(number greater than 3).

OpenStudy (taylor0402):

there can be 6 results if you roll the cube (dice?) once, 1, 2, 3,4,5 ,6 , three numbers 4,5,6 are greater than 3 so P(numbers greater than 3) = n(A)/n(S) = 3/6 = 1/2 I would put 3/6=1/2, put em both down to be safe.

OpenStudy (anonymous):

Thank you (: